# long-var-wrap = dv

### v2

@@ Abstract
Interactions between active particles may be non-reciprocal, breaking action-reaction symmetry and leading to novel physics not seen in equilibrium systems. In this work, we study the state diagram for the non-reciprocal Cahn-Hilliard (NRCH) model with different types of symmetry—discrete and continuous—and examine the effects of non-reciprocity on phase separation. The models are analyzed by drawing the full linear stability diagrams, indicating the onset of phase separation, allowing us to demonstrate how non-reciprocity gives rise to out-of-equilibrium steady states with non-zero currents. These observations remain valid in a large part of parameter space, and with the addition of weak noise. By describing the effects of symmetry on conserved, non-reciprocal models, this work adds to our understanding of the rich behavior of such models.
@@ I Introduction
The physical properties of active matter—comprised of soft matter systems, living or artificial, which continually consume energy—have been the subject of extensive investigations in recent decades [1, 2, 3, 4, 5]. Depending on the mechanisms that underlie the non-equilibrium activity, some of the very basic properties that apply to passive matter will no longer hold for active matter systems. In particular, the action-reaction symmetry exemplified by Newtons third law relies on forces derived from an underlying common interaction potential, which is something one cannot expect in effective descriptions of many active systems such as living organisms. Such non-reciprocal interactions have recently been a topic of interest in active matter, as it is an important feature of effective chemically-mediated [6, 7, 8, 9, 10] or hydrodynamically-mediated [11, 12] interactions, the social interaction of swarms [13], or in models of the brain [14, 15]. Moreover, it has been proposed that naturally occurring non-reciprocal interactions in chemically active systems may have played a role in the emergence of early forms of life [16, 17, 18]. A similar action-reaction asymmetry, exhibited by condensed living matter, is odd elasticity or viscosity [19, 20, 21]. Non-reciprocal interactions are also present in non-active systems driven out of equilibrium, such as driven dusty plasmas [22, 23, 24, 25], magnetic systems [26, 27] or Bose-Einstein condensates [28]. More generally, introducing non-reciprocal interactions is one of the simplest way to break equilibrium, making it useful as a theoretical tool for exploring the behavior non-equilibrium states of matter [29, 13, 30, 31, 32, 33, 34, 35].
@@
Non-reciprocal systems, or more generally non-equilibrium systems, may exhibit exceptional points [36, 37]. These were first considered as features of quantum systems, appearing in models of quantum chaos [38], $\mathcal{PT}$-symmetric systems [39, 40], optical wave-guides [41], and atomic orbitals [42]; all systems where gain and loss result in non-Hermitian terms in the Hamiltonian. In non-reciprocal systems, exceptional points appear in the linearized dynamics of the system, represented by the linear stability matrix [13, 43, 44]. This matrix is in general non-Hermitean due to the broken action-reaction symmetry, and there may therefore be points in parameter space where it becomes defect, its eigenvalues complex, and the corresponding eigenvectors coalesce. Exceptional points in non-equilibrium physics promise to be a fruitful avenue for novel physical effects, such as the recently demonstrated critical fluctuations in driven-dissipative Bose-Einstein condensates [45] or non-equilibrium extensions of $\text{O}(N)$ models [46, 47, 31].
@@
In equilibrium statistical physics, the Cahn-Hilliard model is a widely used phenomenological model for phase separation [48, 49]. The model describes the dynamics of conserved scalar fields $\varphi_{a}(\bm{x},t)$ relaxing to minimize its free energy $F[\varphi]$. The Non-Reciprocal Cahn-Hilliard (NRCH) model extends this model by introducing non-reciprocal interactions that cannot be described in terms of minimizing free energy [43, 44]. This model has recently been shown to exhibit phase behavior beyond that of its well-studied reciprocal counterpart [43, 44, 50, 51, 52].
@@
In this Article, we explore the phase diagram of two restricted versions of the NRCH, one with a $\mathrm{SO}(2)$ continuous rotational symmetry and one with reduced discrete $\mathrm{C}_{4}$-symmetry (see Figure 1). These restricted models allow us to fully characterize their phase diagram at the linear level. We explore how the different characters of these symmetries—continuous and discrete—affect the phase behavior of out-of-equilibrium models. The symmetries are inherited by the linear-stability diagrams, effectively reducing the number of parameters in the continuous case. The symmetry further affects the phase separated state. While, in the discrete case, an energy barrier leads to distinct phases separated by a finite thickness domain wall, there is no barrier in the continuous case, leading to “continuous phase separation”. We further explore how these effects persists even when the system is driven out of equilibrium by non-reciprocal interactions.
@@ II The Model
$\mathcal{F}[\varphi]=\frac{1}{2}\bm{\nabla}\varphi_{a}\cdot\bm{\nabla}\varphi_{a}+\mathcal{V}(\varphi_{a}),$ (1)
where summation is implied. The total free energy reads $F=\int\mathrm{d}\bm{x}\,\mathcal{F}$. The two models are defined by their potential $\mathcal{V}(\varphi)$. The potential of the model with $\mathrm{SO}(2)$ symmetry is
$\mathcal{V}_{\mathrm{cs}}(\varphi_{a})=\frac{r}{2}\varphi_{a}\varphi_{a}+\frac{u}{4}(\varphi_{a}\varphi_{a})^{2}.$ (2)
@@
$\mathcal{V}_{\mathrm{ds}}(\varphi_{a})=\sum_{a}\left[\frac{r}{2}\varphi_{a}^{2}+\frac{u}{2}\varphi_{a}^{4}\right],$ (3)
which is not invariant under continuous rotations, as illustrated in Figure 1. However, $\mathcal{V}_{\mathrm{ds}}$ is invariant under operations in $\mathrm{D}_{4}$, the dihedral group of order 8 corresponding to the symmetries of a square. These operations are rotations with $n\pi/2$ radians in addition to reflections, e.g. $\varphi_{1}\rightarrow-\varphi_{1}$. This is a discrete group, similar to the symmetry group of the single-field Cahn-Hilliard $\mathrm{O}(1)\cong\mathrm{D}_{2}$, and in the same way gives rise to a phase separation with a well-defined domain wall. The introduction of a non-reciprocal term restricts the symmetry to $\mathrm{C}_{4}$, namely, the cyclic group of order 4.
@@
[…]
@@ III Linear Stability
$\partial_{t}\delta\varphi_{a}(\bm{q},t)=-q^{2}\Gamma M_{ab}\delta\varphi_{b}(\bm{q},t).$ (6)
[…]
A non-zero value of $\bar{\varphi}_{a}$ seemingly breaks the assumed symmetry of the model. However, these are still present in the form of rotation of both $\bar{\varphi}$ and $\delta\varphi$. In the case of the continuous symmetry the Ward-Takashi identities related to rotation still apply [55], just as they do when this symmetry is broken spontaneously. The conservation law ensures there are two zero-modes, regardless of whether the continuous symmetry is broken or not, making Goldstone’s theorem non-applicable. In this case, the continuous nature of the symmetry ensures that a domain wall has a divergent length-scale, rather than the correlation length.
@@ III.1 Continuous Symmetry
The base of this diagram is the $\alpha=0$ surface. We choose $\alpha>0$, without loss of generality. At $\alpha=u\bar{\varphi}^{2}$, we have $\xi^{-2}_{+}=\xi^{-2}_{-}$. In reciprocal systems, this indicates a degeneracy where two independent eigenmodes have the same eigenvalue. When we deal with non-reciprocal systems, and thus non-Hermitian matrices, we are no longer guaranteed that the eigenvectors of $M_{ab}$ span the full vector space of perturbations. At this exceptional point (EP) [36, 57], the eigenvalues coincide because the two eigenmodes align, rendering the matrix $M_{ab}$ defective. As there are three free parameters and the exceptional points are given by one equation, namely, $\alpha=u\bar{\varphi}^{2}$, the full set of exceptional points forms a 2D surface, which we denote as the exceptional plane. This is illustrated in purple in the Figure 2. Above the exceptional plane, the two eigenvalues become complex, and the perturbations will no longer decay or diverge monotonically: they exhibit oscillatory behavior. Here, we have $\mathrm{Re}\left(\xi_{\pm}^{-2}\right)=r+2u\bar{\varphi}^{2}$, and, therefore, both modes transition from stable to unstable at the same surface, defined via $r=-2u\bar{\varphi}^{2}$. The intersection between the critical and the exceptional planes can be called critical exceptional points (CEPs) [45]. In the dynamical systems literature, these are know as Takens-Bogdanov bifurcations [58]. These CEPs are parameterized as follows
@@
Figure 3: The four non-trivial slices of the full three-dimensional state diagram shown in Figure 2. Panel (a) corresponds to the equilibrium case $\alpha=0$, while panels (b)-(d) illustrate three orthogonal slices for finite $\alpha$, at constant $\alpha$, $\bar{\varphi}$, and $r$, respectively. Light blue and hashed areas mark the type-A and type-B unstable regions. Black, blue dash-dotted, and green lines separate stable and type-A, type-A and type-B, and stable and type-B, respectively. The purple dashed line marks the exceptional points, and the red dot marks the critical exceptional points where all the lines intersect.
@@
Looking back at the Mexican hat potential illustrated in Figure 1, this corresponds to a point inside the ring of minima, and perturbations that would correspond to the Goldstone mode in a spontaneous symmetry breaking scenario (i.e. the orthogonal to the radial direction). The mode corresponding to perturbations in the radial direction ($\delta\varphi_{1}$ in this case) remains stable. This holds in the light blue area labeled “type-A unstable”, where there is one unstable mode. At $u\bar{\varphi}^{2}=|r|/3$, illustrated by a blue dash-dotted line, the second mode also transitions from stable to unstable. The area where both modes are unstable is highlighted as dark blue and hashed, with the label “type-B unstable”.
@@
$\frac{r}{\alpha}=-2\,\frac{u\bar{\varphi}^{2}}{\alpha}\pm\sqrt{\frac{u^{2}\bar{\varphi}^{4}}{\alpha^{2}}-1},\quad\sqrt{u}|\bar{\varphi}|>\alpha.$ (10)
@@
$r=-2u\bar{\varphi}^{2},\quad u|\bar{\varphi}|\leq\alpha,$ (11)
@@ III.2 Discrete Symmetry
Since the eigenvalues depend on both $\bar{\varphi}_{1}^{2}$ and $\bar{\varphi}_{2}^{2}$, the full linear stability diagram requires another dimension to visualize. For this model, we focus on a 3D slice of the full 4D diagram as shown in Figure 4. This cross section is related to any other 3D slice via a choice of scaling, which in this case is chosen to be $|r|$. The inverse-squared correlation lengths defined via $\xi_{\pm}^{-2}=\lambda_{\pm}(q\rightarrow 0)$ as above are then given as follows
$\xi_{\pm}^{-2}$ $=$ $-|r|+3u\left(\bar{\varphi}_{1}^{2}+\bar{\varphi}_{2}^{2}\right)$ (16)
$\pm$ $\sqrt{9u^{2}\left({\bar{\varphi}_{1}^{2}}-{\bar{\varphi}_{2}^{2}}\right)^{2}-\alpha^{2}}.$
Figure 5: The critical line for $\alpha/|r|\in[0,2]$. Brighter colors in the plot indicate higher values of $\alpha$.
Figure 6: Constant $\alpha$ slices of the linear stability diagram illustrated in Figure 4. The gray dashed lines represent the line connecting the minima of the potential illustrated in Figure 1(a), and form a square centered at the origin and with side lengths $\sqrt{u/|r|}\bar{\varphi}=\sqrt{2}$. Light blue and hashed areas mark the type-A and type-B unstable regions. Black, blue dash-dotted, and green lines separate stable and type-A, type-A and type-B, and stable and type-B, respectively. The purple line marks the exceptional points, and the red dot marks the critical exceptional points where all the lines intersect.
@@
with a similar solution for $\bar{\varphi}_{1}$. This sector, which is plotted for $u\bar{\varphi}_{1}^{2}/|r|\in[0,1/\sqrt{6}]$ in Figure 5, represents one eighth of the total solution set, with the rest being obtained by applying the operations of $\mathrm{D}_{4}$. The critical surface is highlighted in green in Figure 4. For a comparison with the continuous case, Figure 4 corresponds to Figure 3(d).
For $\alpha=0$ the critical lines are straight, where $\sqrt{u/|r|}\bar{\varphi}_{1}$ or $\sqrt{u/|r|}\bar{\varphi}_{2}$ are ±$\pm\frac{1}{\sqrt{6}}\approx\pm 0.4$. As $\alpha$ is increased, the critical lines become curved, asymptotically approaching the $\alpha=0$ lines for large values of the fields. The exceptional plane, shown in purple in Figure 4, is given by a set of hyperbolae defined via $u|\bar{\varphi}_{1}^{2}-\bar{\varphi}_{2}^{2}|=\alpha/3$, where we again have chosen $\alpha>0$ for concreteness. Inside the area bounded by the exceptional surface, i.e. the subset containing the origin, the critical surface is given by the circle $u\left(\bar{\varphi}_{1}^{2}+\bar{\varphi}_{2}^{2}\right)=|r|/3$ where $\mathrm{Re}\left(\xi_{\pm}^{-2}\right)=0$, which is shown in green in Figure 4.
The intersection of the critical surface (highlighted as green in Figure 4) and the exceptional surface (highlighted as purple in Figure 4) is the critical exceptional line, which is shown in red in Figure 4 [see the corresponding equivalent plot for the continuous symmetry case, in Figure 3(d)]. These are located at the intersections of the circle of radius $1/\sqrt{3}$ and the hyperbolas with semi-major axis $\sqrt{\alpha/(3|r|)}$. For $\alpha>|r|$ there are no such points, for $\alpha=0$ and $\alpha=|r|$ there are four, while for $\alpha/|r|\in(0,1)$ there are eight, corresponding to the elements of $\mathrm{D_{4}}$. These observations can be seen in the two-dimensional slices of the linear stability diagram shown in Figure 6. Additional variations of such plots are presented in the SM[56].

### v3

# State Diagram of the Non-Reciprocal Cahn-Hilliard Model and the Effects of Symmetry

@@ Abstract
Interactions between active particles may be non-reciprocal, breaking action-reaction symmetry and leading to novel physics not observed in equilibrium systems. The non-reciprocal Cahn-Hilliard (NRCH) model is a phenomenological model that captures the large-scale effects of non-reciprocity in conserved, phase-separating systems. In this work, we explore the consequences of different variations of this model corresponding to different symmetries, inspired by the importance of symmetry in equilibrium universality classes. In particular, we contrast two models, one with a continuous $\mathrm{SO}(2)$ symmetry and one with a discrete $\mathrm{C}_{4}$ symmetry. We analyze the corresponding models by constructing three-dimensional linear stability diagrams. With this, we connect the models with their equilibrium limits, highlight the role of mean composition, and classify qualitatively different instabilities. We further demonstrate how non-reciprocity gives rise to out-of-equilibrium steady states with non-zero currents and present representative closed-form solutions that help us understand characteristic features of the models in different parts of the parameter space.
@@ I Introduction
The physical properties of active matter—comprising soft matter systems, living or artificial, which continually consume energy—have been the subject of extensive investigations in recent decades [1, 2, 3, 4, 5]. Depending on the mechanisms that underlie the non-equilibrium activity, some of the very basic properties that apply to passive matter will no longer hold. In particular, the action-reaction symmetry exemplified by Newtons third law relies on forces derived from an underlying common interaction potential, something one cannot expect in effective descriptions of many active systems such as living organisms. Such non-reciprocal interactions have recently been a topic of interest in active matter, as it is an important feature of effective chemically-mediated [6, 7, 8, 9, 10] or hydrodynamically-mediated [11, 12] interactions, the social interaction of swarms [13], or in models of the brain [14, 15]. Moreover, it has been proposed that naturally occurring non-reciprocal interactions in chemically active systems may have played a role in the emergence of early forms of life [16, 17, 18]. A similar action-reaction asymmetry, exhibited by condensed living matter, is odd elasticity or viscosity [19, 20, 21]. Non-reciprocal interactions are also present in non-active systems driven out of equilibrium, such as driven dusty plasmas [22, 23, 24, 25], magnetic systems [26, 27] or Bose-Einstein condensates [28]. More generally, introducing non-reciprocal interactions is one of the simplest way to break equilibrium, making it useful as a theoretical tool for exploring the behavior non-equilibrium states of matter [29, 13, 30, 31, 32, 33, 34, 35].
@@
Non-reciprocal systems, or more generally non-equilibrium systems, may exhibit exceptional points [36, 37]. These were first considered as features of quantum systems, appearing in models of quantum chaos [38], $\mathcal{PT}$-symmetric systems [39, 40], optical wave-guides [41], and atomic orbitals [42]; all systems where gain and loss result in non-Hermitian terms in the Hamiltonian. In non-reciprocal systems, exceptional points appear in the linearized dynamics of the system, represented by the linear stability matrix [13, 43, 44]. This matrix is in general non-Hermitian—i.e., it does not equal its transposed complex-conjugate—due to the broken action-reaction symmetry, and there may therefore be points in parameter space where it becomes defective, its eigenvalues complex, and the corresponding eigenvectors coalesce [^1] Since the inner product in this case is strictly $L^{2}$ (including not only sum over components but also integral over the domain), the Hermitian-conjugate also includes integration by parts.. Exceptional points in non-equilibrium physics promise to be a fruitful avenue for novel physical effects, such as the recently demonstrated critical fluctuations in driven-dissipative Bose-Einstein condensates [45] or non-equilibrium extensions of $\text{O}(N)$ models [46, 47, 31].
@@
In equilibrium statistical physics, the Cahn-Hilliard model is a widely used phenomenological model for phase separation [48, 49]. The model describes the dynamics of conserved scalar fields $\varphi_{a}(\bm{x},t)$ relaxing to minimize its free energy $F[\varphi]$. The Non-Reciprocal Cahn-Hilliard (NRCH) model extends this model by introducing non-reciprocal interactions that cannot be described in terms of minimizing free energy [43, 44]. This model has recently been shown to exhibit phase behavior beyond that of its well-studied reciprocal counterpart [43, 44, 50, 51, 52]. As a phenomenological model, NRCH has a priori many independent parameters. In the literature, many different versions of NRCH have been studied based on the specific focus of the intended investigation, such as a version where the two species are treated on an equal footing [43], a version where one species has minimal dynamics [44], and a version where rotational invariance in the space of species is introduced [43, 51, 52, 53]. In this paper, we aim to investigate the role of symmetry with respect to transformations in the space of fields on the behavior of the NRCH model, and take the first steps towards categorizing the phenomenology of its different variations.
@@
To explore the consequences of symmetry, we consider two versions of the model: one with the discrete $\mathrm{C}_{4}$-symmetry and one with the continuous $\mathrm{SO}(2)$ symmetry (see Figure 1). These restrictions allow us to fully characterize their state diagram at the linear level. The symmetries are inherited by the linear-stability diagrams, reducing the number of free parameters in the continuous case. In the discrete case, an energy barrier leads to distinct phases separated by a finite thickness domain wall, while in the continuous case there is no barrier, leading to “continuous phase separation”. We further explore how these effects persists even when the system is driven out of equilibrium by non-reciprocal interactions.
@@ II The Model
$\mathcal{F}[\varphi]=\frac{1}{2}\bm{\nabla}\varphi_{a}\cdot\bm{\nabla}\varphi_{a}+\mathcal{V}(\{\varphi_{a}\}),$ (1)
where summation over repeated indices is implied. The total free energy reads $F=\int\mathrm{d}\bm{x}\,\mathcal{F}$. The two models are defined by their potential $\mathcal{V}(\varphi)$. The potential of the model with $\mathrm{SO}(2)$ symmetry is
$\mathcal{V}_{\mathrm{cs}}(\{\varphi_{a}\})=\frac{r}{2}\varphi_{a}\varphi_{a}+\frac{u}{4}(\varphi_{a}\varphi_{a})^{2}.$ (2)
@@
$\mathcal{V}_{\mathrm{ds}}(\{\varphi_{a}\})=\sum_{a}\left[\frac{r}{2}\varphi_{a}^{2}+\frac{u}{2}\varphi_{a}^{4}\right],$ (3)
which is not invariant under continuous rotations, as illustrated in Figure 1. However, $\mathcal{V}_{\mathrm{ds}}$ is invariant under operations in $\mathrm{D}_{4}$, the dihedral group of order 8 corresponding to the symmetries of a square. These operations are rotations with $n\pi/2$ radians in addition to reflections, e.g. $\varphi_{1}\rightarrow-\varphi_{1}$. This, similarly to the the symmetry group of the single-field Cahn-Hilliard $\mathrm{O}(1)\cong\mathrm{D}_{2}]$, is a discrete group and therefore gives rise to a phase separation with a well-defined domain wall. The introduction of a non-reciprocal term restricts the symmetry to $\mathrm{C}_{4}$, namely, the cyclic group of order 4.
@@
$\partial_{t}\varphi_{a}=\Gamma\nabla^{2}\left[{\frac{\partial\mskip 0.0mu\mathcal{V}}{\partial\varphi_{a}}}-\nabla^{2}\varphi_{a}+\alpha\epsilon_{ab}\varphi_{b}\right].$ (5)
@@ III Linear Stability
$\partial_{t}\delta\varphi_{a}(\bm{q},t)=-q^{2}\Gamma M_{ab}(\bm{q})\delta\varphi_{b}(\bm{q},t).$ (6)
The matrix $M_{ab}$ depends on the specific choice for the potential, and thus reflects the effects of the underlying symmetry. We set $\Gamma=1$, which can be done without loss of generality by scaling time $t$.
A non-zero value of $\bar{\varphi}_{a}$ seemingly breaks the assumed symmetry of the model. However, the symmetry is still present in the form of rotation of both $\bar{\varphi}$ and $\delta\varphi$. In the case of the continuous symmetry the Ward-Takahashi identities related to rotation still apply [55], just as they do when this symmetry is broken spontaneously. It might appear tempting to invoke Goldstone’s theorem in this case, treating $\bar{\varphi}$ as the order-parameter whose non-zero value may indicate a broken symmetry giving rise to slow modes. However, as the system already features two slow-modes due to the conservation laws, the breaking of the continuous symmetry of $\varphi$-rotation does not yield any additional slow-modes, and thus Goldstone’s theorem does not provide any information about the dynamics of the system [^2] In terms of the vertex-functions $\Gamma_{ab}(\omega,\bm{q})$—the inverse Greens-functions—Goldstone’s theorem states that $\left\langle\varphi_{a}\right\rangle\Gamma_{ab}(\bm{q}=\omega=0)=0$ [56]. So, either the symmetry is preserved ($\left\langle\varphi_{a}\right\rangle=\bar{\varphi}_{a}=0$), or $\Gamma$ has a zero. However, the conservation laws ensure that $\Gamma(\omega=0)\propto q^{2}$, independently of whether or not the symmetry is broken.. In fact, the value of $\bar{\varphi}_{a}$ is a fixed parameter, and not chosen spontaneously by the system.
@@ III.1 Continuous Symmetry
Figure 3: The four non-trivial slices of the full three-dimensional state diagram shown in Figure 2. Panel (a) corresponds to the equilibrium case $\alpha=0$, while panels (b)-(d) illustrate three orthogonal slices for finite $\alpha$, at constant $\alpha$, $\bar{\varphi}$, and $r$, respectively. Light blue and hashed areas mark the type-A and type-B unstable regions. Black, blue dash-dotted, and green lines separate stable and type-A, type-A and type-B, and stable and type-B, respectively. The purple dashed line marks the exceptional points, the yellow dot the critical point, and the red dots the critical exceptional points where all the lines intersect.
@@
The base of this diagram is the $\alpha=0$ surface. We choose $\alpha>0$, without loss of generality. At $\alpha=u\bar{\varphi}^{2}$, we have $\xi^{-2}_{+}=\xi^{-2}_{-}$. In reciprocal systems, this indicates a degeneracy where two independent eigenmodes have the same eigenvalue. When we deal with non-reciprocal systems, and thus non-Hermitian matrices, we are no longer guaranteed that the eigenvectors of $M_{ab}$ span the full vector space of perturbations. At this exceptional point (EP) [36, 58], the eigenvalues coincide because the two eigenmodes align, rendering the matrix $M_{ab}$ defective. As there are three free parameters and the exceptional points are given by one equation, namely, $\alpha=u\bar{\varphi}^{2}$, the full set of exceptional points forms a 2D surface, which we denote as the exceptional plane. This is illustrated in purple in the Figure 2. Above the exceptional plane, the two eigenvalues become complex, and the perturbations will no longer decay or diverge monotonically: they exhibit oscillatory behavior. Here, we have $\mathrm{Re}\left(\xi_{\pm}^{-2}\right)=r+2u\bar{\varphi}^{2}$, and, therefore, both modes transition from stable to unstable at the same surface, defined via $r=-2u\bar{\varphi}^{2}$. The intersection between the critical and the exceptional planes can be called critical exceptional points (CEPs) [45]. In the dynamical systems literature, these are known as Takens-Bogdanov bifurcations [59]. These CEPs are parameterized as follows
@@
This diagram is similar to the binodal-spinodal diagram for phase separation of one species. In that case, the outer line represents the binodal, within which the the system is metastable, while the inner line gives the spinodal, where the system becomes linearly unstable. In this case, on the other hand, the system has an additional soft direction (orthogonal to the radial direction) as a result of the rotational symmetry. The outer line (the type-A border) therefore represents the onset of linear instability, while also being the binodal, and no metastable state exists. The inner line (the type-B border) is where the second eigenvalue changes sign. At $r=\bar{\varphi}=\alpha=0$ is the well-known $\mathrm{O}(2)$ critical point. For $\alpha\neq 0$, the outer line denotes linear instabilities, but the identifying binodals becomes more complicated [60, 61]. For small $\alpha$, we expect the instabilities to be a good approximation for the binodals.
Looking back at the Mexican hat potential illustrated in Figure 1, this corresponds to a point inside the ring of minima, and perturbations that would correspond to the Goldstone mode in a spontaneous symmetry breaking scenario (orthogonal to the radial direction). The mode corresponding to perturbations in the radial direction, $\delta\varphi_{1}$, remains stable. This holds in the light blue area labeled “type-A unstable”, where there is one unstable mode. At $u\bar{\varphi}^{2}=|r|/3$, illustrated by a blue dash-dotted line, the second mode also transitions from stable to unstable. The area where both modes are unstable is highlighted as dark blue and hashed, with the label “type-B unstable”.
@@
$\frac{r}{\alpha}$ $=-2\,\frac{u\bar{\varphi}^{2}}{\alpha}\pm\sqrt{\frac{u^{2}\bar{\varphi}^{4}}{\alpha^{2}}-1},$ $u|\bar{\varphi}|^{2}$ $>\alpha.$ (10)
@@
$r$ $=-2u\bar{\varphi}^{2},$ $u|\bar{\varphi}|^{2}$ $\leq\alpha,$ (11)
@@ III.2 Discrete Symmetry
Since the eigenvalues depend on both $\bar{\varphi}_{1}^{2}$ and $\bar{\varphi}_{2}^{2}$, the full linear stability diagram requires another dimension to visualize. For this model, we focus on a 3D slice of the full 4D diagram as shown in Figure 4. This cross section is related to any other 3D slice via a choice of scaling, which in this case is chosen to be $|r|$. 2D cross sections of Figure 4 for $\alpha=\mathrm{const.}$ is shown in Figure 5. For a comparison with the continuous case, Figure 4 corresponds to Figure 3(d). The inverse-squared correlation lengths defined via $\xi_{\pm}^{-2}=\lambda_{\pm}(q\rightarrow 0)$ as above are then given as follows
$\xi_{\pm}^{-2}$ $=-|r|+3u\left(\bar{\varphi}_{1}^{2}+\bar{\varphi}_{2}^{2}\right)\pm\sqrt{9u^{2}\left({\bar{\varphi}_{1}^{2}}-{\bar{\varphi}_{2}^{2}}\right)^{2}-\alpha^{2}}.$
@@
with a similar solution for $\bar{\varphi}_{1}$. See the Supplementary Material [57] for plots of this for a range of parameter values. The critical surface is highlighted in green in Figure 4.
Figure 5: Constant $\alpha$ slices of the linear stability diagram illustrated in Figure 4. The gray dashed lines represent the line connecting the minima of the potential illustrated in Figure 1(a), and form a square centered at the origin and with side lengths $\sqrt{u/|r|}\bar{\varphi}=\sqrt{2}$. Light blue and hashed areas mark the type-A and type-B unstable regions. Black, blue dash-dotted, and green lines separate stable and type-A, type-A and type-B, and stable and type-B, respectively. The purple lines mark the exceptional points, the yellow dots the critical point, and the reds dot the critical exceptional points where all the lines intersect.
For $\alpha=0$ the critical lines are straight, where $\sqrt{u/|r|}\bar{\varphi}_{1}$ or $\sqrt{u/|r|}\bar{\varphi}_{2}$ are $\pm 1/\sqrt{6}\approx\pm 0.4$. As $\alpha$ is increased, the critical lines become curved, asymptotically approaching the $\alpha=0$ lines for large values of the fields. The exceptional plane, shown in purple in Figure 4, is given by a set of hyperbolae defined via $u|\bar{\varphi}_{1}^{2}-\bar{\varphi}_{2}^{2}|=\alpha/3$, where we again have chosen $\alpha>0$ for concreteness. Inside the area bounded by the exceptional surface, i.e., the subset containing the origin, the critical surface is given by the circle $u(\bar{\varphi}_{1}^{2}+\bar{\varphi}_{2}^{2})=|r|/3$ where $\mathrm{Re}(\xi_{\pm}^{-2})=0$, which is shown in green in Figure 4. The intersection of the critical surface and the exceptional surface is the critical exceptional line, which is shown in red in Figure 4. These are located at the intersections of the circle of radius $1/\sqrt{3}$ and the hyperbolas with semi-major axis $\smash{\sqrt{\alpha/(3|r|)}}$. For $\alpha>|r|$ there are no such points, for $\alpha=0$ and $\alpha=|r|$ there are four, while for $\alpha/|r|\in(0,1)$ there are eight, corresponding to the elements of $\mathrm{D_{4}}$. These observations can be seen in the two-dimensional slices of the linear stability diagram shown in Figure 5. Additional variations of such plots are presented in the supplemental material [57].